A positive number is divided into two parts such that the squares of the two parts is 20. The square of the larger number is 8 times the smaller number. Take the smaller number as x.

Answer by Theo(13342)

x is the smaller number.
y is the larger number.
x^2 + y^2 = 20
you are given that the square of the larger number is equal to 8 time the smaller number.
that gets you y^2 = 8x.
rpelace y^2 with 8x in the original equation to get:
x^2 + 8x = 20
subtract 20 from both sides of the equation to get:
x^2 + 8x - 20 = 0
factor this quadratic equation to get:
(x+10) * (x-2) = 0
solve for x to get:
x = -10 and x = 2
since x has to be positive, you get x = 2 as your potential solution.
when x = 2, y^2 = 8x = 16.
this makes y = 4
your solution is that x = 2 and y = 4
x^2 + y^2 = 20 becomes 2^2 + 4^2 = 20 which becomes 4 + 16 = 20 which becomes 20 = 20.
this confirms the values are good.
you get:
x = 2 and y = 4 as your solution.